Appendix A. ACTION-NYS

This project is part of our new industrial economic development program ACTION-NYS, which is partially funded by New York State. We hope that this project will be one of our showcase industrial applications. The State funding will allow us to reduce the level of funding from Niagara Mohawk, since ACTION will cover a substantial part of the costs of the proposal.

ACTION-NYS is the technology transfer program of the Northeast Parallel Architectures Center (NPAC). It introduces and integrates parallel computing into New York State industry by enabling NPAC scientists and affiliated Syracuse faculty to apply their parallel computing knowledge to large, complex industrial applications. ACTION also provides industrial access to the latest parallel machines, as well as education and training in high performance parallel computing.

Direction for the ACTION program is provided by Professor Geoffrey Fox with guidance from the ACTION High Performance Computing Industrial Advisory Board consisting of 17 top executives from New York State industries and the ACTION Regional Economic Development Committee comprised of representatives from Central New York economic development agencies. Niagara Mohawk is represented on the Advisory Board.

The ACTION Program operates by reaching out to New York State industry to explore opportunities in parallel computing. In cooperation with high level technical executives of targeted corporations, ACTION initially prepares a survey of potential for parallel computing for that industry. The experienced NPAC staff then determines the most effective methodology for integration of parallel computing into the industry. Typically, small “proof-of-concept” projects are initiated to demonstrate that parallel computing is beneficial to the industry. The next step is to identify a large applications project that actually begins to integrate parallel computing technology into the processes of the corporation. A number of industrial projects are already completed or underway, with such companies as IBM, General Electric, and Sonnet Software.

Other activities of ACTION include the quarterly publication of the newsletter, Parallel Computing News, produced to inform NPAC’s clients and advisors of current developments in the field of high performance parallel computing. Education and training are also offered to prepare the next generation of workers and to upgrade skills of the present workforce. Geoffrey Fox and other ACTION-affiliated experts often address groups of industrial executives to introduce them to issues and uses of parallel computing.

High performance parallel computing can play an important role in the development of industries and the advancement of technologies that lead to industrial development. The ACTION program contributes to this effort through the dissemination, application and integration of parallel computing to help industries to develop new consumer and industrial products, and the development of new processes to increase productivity and the competitiveness of existing products.